Please rate I’ll give a thumbs up!
Consider the relationship between the number of bids an item on eBay received and the items selling price. The following is a sample of 5 items sold through an auction. Price in Dollars 29 32 40 45 49 Number of Bids 23 7 89 Table Copy Data Step 1 of S: Calculate the sum of squared errors (SSE). Use the values bo-8.3804 and b10.3636 for the calculations. Round your answer to three decimal places. AnswerHow to Enter) 2 Points Keypad Tables rev Next
tep 2 of S: Calculate the estimated variance of errors, s. Round your answer to three decimal places.
ep 3 of S: Calculate the estimated variance of slope, s. Round your answer to three decimal places.
Step 4 of 5: Construct the 95 % confidence interval for the slope. Round your answers to three decimal places. Answer(How to Enter) 4 Points Keypad Tables Lower endpoint: Nex Upper endpoint
Step 5 of 5: Construct the 90 % confidence interval for the slope. Round your answers to three decimal places. Answer(How to Enter) 2 Points I Keypad m Tables ! Lower endpoint Upper endpoint Nek
Answer #1

data

price in dollar number of bids
29 2
32 3
40 7
45 8
49 9

Using Excel

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.987266635
R Square 0.974695408
Adjusted R Square 0.966260544
Standard Error 0.572077554
Observations 5
ANOVA
df SS MS F Significance F
Regression 1 37.81818182 37.81818182 115.5555556 0.001721541
Residual 3 0.981818182 0.327272727
Total 4 38.8
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 90.0% Upper 90.0%
Intercept -8.381818182 1.343856545 -6.237137598 0.008312219 -12.65856948 -4.105066886 -11.54440104 -5.219235327
price in dollar 0.363636364 0.033827655 10.749677 0.001721541 0.255981668 0.471291059 0.284027598 0.44324513

1)

SSE = 0.982

2)

se^2 = MSE = 0.327

3)

s^2_b = 0.0338^2 = 0.0011

4)

95% confidence

lower = 0.256

upper = 0.471

5)

90% confidence level

lower = 0.284

upper = 0.443